OpenStudy (anonymous):

-2y-4=8-y can someone show me step by step because im messing up somewhere

OpenStudy (anonymous):

@herobeautiful -2y -4 = 8 -y simplify both sides of the equation -2y -4 = -y + 8 add y to both of the sides -2y - 4 + y = -y + 8 + y -y -4 = 8 add 4 both sides -y -4 + 4 = 8 + 4 -y = 12 divide both sides by -1 -y/-1 = 12/-1 answer -> y = -12